HUBBLE SPACE TELESCOPE DAILY REPORT #2458 PERIOD COVERED: 0000Z (UTC) 09/15/99 - 0000Z (UTC) 09/16/99 Daily Status Report as of 259/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 8415 (CCD Sparse-Field CTE External C8) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of NGC-6752, a cluster associated with a nebula in Tucana. This program will allow the derivation of accurate coefficients that can be used to correct low count level data {both imaging and spectroscopy} for nonlinearity/CTE effects. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 8443 (WF/PC-2 Cycle-8 Supplemental Darks pt1/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no problems were reported. 1.3 Completed Eight Sets of WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 8301 (A Snapshot Survey of X-ray Selected Central Cluster Galaxies) The WF/PC-2 was used to make snapshot survey observations of the x-ray, central cluster galaxy A697. Central cluster galaxies are the most massive stellar systems known and have been used as standard candles for many decades. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Seven Sets of STIS/CCD 8406 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation V) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations taken in parallel with the prime science observations as part of a POMS proposal. This is a STIS non-scripted parallel Condition V proposal. The observations were completed as scheduled, and no anomalies were reported. 1.6 Completed STIS/CCD 8408 (STIS CCD Dark Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the darks for the STIS CCD. The observations were executed as scheduled, and no anomalies were reported. 1.7 Completed Three Sets of WF/PC-2 8194 (Evolution of High-Redshift Seyfert Galaxies) The WF/PC-2 was used to make observations of the high-redshifted Seyfert galaxies Gal-105120, Gal-105312 and Gal-105108in order to study their evolution. The observations will provide insights into the Seyfert phenomenon at intermediate redshift and will complement similar HST studies of normal galaxies, clusters and quasars. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed STIS/CCD 8409 (STIS CCD Bias Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and to track the evolution of hot columns. The observations were executed as scheduled, and no anomalies were reported. 1.9 Completed FGS 8463 (Long Term Stability of FGS1R in Transfer Scan Mode) The Fine guidance Sensors were used to make astrometric observations of star UPGREN69 and several other reference stars in order to study the long term stability of FGS-1R in the Transfer Scan Mode. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed STIS/CCD 8172 (STIS Deep DLA Imaging Survey II) The Space Telescope Imaging Spectrograph (CCD) was used to make deep observations of quasar/alpha galaxy Q2233+131. The HST imaging, together with ground-based spectroscopy, will unify absorption-line studies and faint galaxy imaging studies, providing a comprehensive picture of the history of assembly, star formation, gas consumption, and chemical enrichment in the Damped Lyman-Alpha galaxies. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed STIS/CCD 8313 (Imaging the Host Galaxies of High Redshift Type Ia Supernovae) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of 94102 and its host galaxy. The propose is to obtain snapshot, unfiltered STIS images of distant galaxies of known redshift which have hosted supernovae {SNe} of Type IA found via the Supernova Cosmology Project. The observations were completed as scheduled, and no anomalies were reported. 1.12 Completed WF/PC-2 8436 (POMS Test Proposal: WFII Backup Parallel Archive Proposal II) The WF/PC-2 was used to make archive observations parallel to the prime science observations as part of a POMS test proposal. This is a POMS test proposal designed to simulate scientific plans. The observations were completed as planned, and no anomalies were reported. 1.13 Completed STIS/CCD/MA1 8280 (Origins, Structure and Evolution of Magnetic Activity in the Cool Half of the H-R Diagram: A STIS Survey) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make observations of star HD22468 to carry out part of a major ultraviolet spectral survey of "coronal" late-type stars. A wavelength calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed Four Sets of WF/PC-2 8457 (UV Earthflats) The WF/PC-2 was used to make a ultraviolet Earth flat calibration image. This is to improve the quality of pipeline the ultraviolet flatfields. The observations were completed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 8 Successful: 8 Scheduled Re-acquisitions: 4 Successful: 4 2.2 FHST Updates: Scheduled: 19 Successful: 19 2.3 Operations Notes: The SSR EDAC error counter was cleared three times per ROP SR-1. The NICMOS FPA temperature sensor for detector 2 was enables at 258/1044Z. Using ROP SD-4, an SSR engineering record was commanded at 258/1535Z. The secondary mirror was moved from 2.7 microns to +0.5 microns, a total move of +3.2 microns, in five steps. The move was completed by 258/1551Z. A TRR was written when GCMRs could not be issued due to an NCC problem. This started at 258/1758Z. Per HSTAR 7255, the voltages for batteries 4, 3, 2, and 1 flagged low at 27.64 volts (the low limit is 27.64 volts) several times starting at 258/181747Z. This was near the end of an orbit night period. Real-time operations switched from CCS “B” string to CCS “A” string at 258/2000Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH